# Unravelling the emergence of quantum jumps in a monitored qubit

### v1

@@ Abstract
To probe this transition, we perform state tomography after variable evolution times without conditioning on the click record (Fig. 5a). For small $\lambda$, the decay rate grows monotonically with $\lambda$, while the oscillation frequency decreases (Fig. 5b). However, at the observed transition point $\lambda^{\mathrm{obs}}_{3}=1.09\pm 0.01$ (Methods), the dynamics become overdamped and oscillations vanish. Beyond this point, further increasing $\lambda$ slows down the exponential decay, completing the entry into the quantum Zeno regime [38]. The observed transition lies well below the ideal prediction $\lambda^{\mathrm{c}}_{3}=2$ owing to the finite waiting time $\tau_{\mathrm{B}}$ in the detector’s excited state between a click and the start of a new trajectory. This waiting time renders the dynamics partially non-Hermitian, lowering the critical value (Supplementary Information Section D[22].
Figure 6: Summary of transition locations. Comparison of the three critical measurement strengths $\lambda^{c}_{1,2,3}$ across three levels of description: the ideal model ($\lambda^{\mathrm{c}}_{1}=1$, $\lambda^{\mathrm{c}}_{2}=2/\sqrt{3}$, $\lambda^{\mathrm{c}}_{3}=2$), the realistic model (Supplementary Information Section 5) incorporating decoherence ($\lambda^{\mathrm{c}}_{1}=1.18$, $\lambda^{\mathrm{c}}_{2}=1.01$, $\lambda^{\mathrm{c}}_{3}=1.25$), and experimental observations ($\lambda^{\rm obs}_{1}=0.99$, $\lambda^{\rm obs}_{2}=0.92$, $\lambda^{\rm obs}_{3}=1.09$). Notably, decoherence inverts the ordering of the first two transitions ($\lambda^{\mathrm{c}}_{2}<\lambda^{\mathrm{c}}_{1}$), a feature captured by the realistic model. Remaining discrepancies likely reflect residual systematics in calibration and data analysis (Supplementary Information Section 7).
@@ Conclusions
The transition from drive-dominated coherent evolution to measurement-dominated behaviour manifests not as a gradual crossover, but as a discrete cascade of dynamical phases. By accessing the full measurement record, we resolve these transitions individually, each revealing a distinct hallmark of monitored quantum systems that is obscured in ensemble averages. Strikingly, decoherence does not blur these transitions but fundamentally restructures the dynamical phase diagram (Fig. 6), inverting the ordering of the first two critical points: dephasing shifts the state-freezing threshold below the onset of continuous quantum jumps, while the detector’s finite waiting time lowers the onset of quantum Zeno dynamics (Supplementary Information Section 5). Additional probes—such as click counting fields [18] or trajectory-resolved entropy production [13]—may uncover further structure, including topologically distinct phases [18, 30]. More broadly, the cascaded emergence of measurement-induced dynamics should extend beyond the single-qubit setting, with implications for measurement-enhanced entanglement generation [20] and measurement-induced phase transitions in many-body systems [36, 19, 39, 10].
@@ Acknowledgments
B.G. and D.G. designed and performed the experiments. B.G., D.G., and N.N. analysed data and performed numerical simulations. L.M.J. performed the device fabrication. U.G. and S.H. developed code for the experiments. P.K., A.R., K.S., and Y.G. developed the theory. B.G., D.G., N.N., K.S., A.R., P.K., Y.G., and S.R. wrote the manuscript, with input from all authors. Y.G., F.L., K.S., and S.R. conceived and supervised the project.
@@
SI 1 Experimental System
@@
Figure S1: Experimental implementation of a qubit with tunable monitoring strength. a, Two-transmon chip mounted within a rectangular 3D superconducting cavity. In the expanded view, the detector (“bright”) transmon (red) is aligned with the cavity field lines (yellow arrows) to ensure strong dispersive coupling. The system (“dark”) transmon (blue) is oriented perpendicular to the field, strongly suppressing its dispersive coupling to the readout mode. b, Energy level diagram showing the shared ground state $\left|0\right\rangle$ and two single-excitation states $\left|1\right\rangle$ and $\left|\mathrm{B}\right\rangle$ forming a V-type structure. Higher-excitation states (e.g., the doubly excited state, shown in grey) lie outside the operating manifold. A resonant drive of amplitude $\Omega_{\mathrm{B}}$ couples $\left|0\right\rangle$ to $\left|\mathrm{B}\right\rangle$, whose population is continuously monitored via the readout cavity at rate $\Gamma$. c, This configuration effectively realises a two-level system whose ground state is continuously monitored by a click detector with tunable measurement strength $\alpha\propto\Omega_{\mathrm{B}}^{2}/\Gamma$.
SI 2 System Parameters
@@ A Drive Configuration
To maximise detection fidelity for $\left|\mathrm{B}\right\rangle$ while minimising decoherence in the system manifold $\{\left|0\right\rangle,\left|1\right\rangle\}$, we drive the readout at frequency $\omega_{\mathrm{R}}-\chi_{\mathrm{B}}$, which is resonant with the cavity only when the system is in $\left|\mathrm{B}\right\rangle$. The resulting monitoring rate of the $\left|\mathrm{B}\right\rangle$ population is [4]
@@ B Continuous Monitoring Protocol
Figure S2: Click record extraction. a, Raw integrated measurement signal versus time. b, Histogram of measurement outcomes, fitted with a sum of two Gaussians, with typical false-negative/false-positive probabilities of $p_{\mathrm{FP}}=p_{\mathrm{FN}}=4\%$. The dashed line indicates the discrimination threshold; the grey band marks the uncertainty region used for Schmitt-trigger filtering. c, Resulting binary click record. d, Distribution of no-click durations with $\Omega_{\mathrm{S}}=0$, showing exponential decay at rate $\alpha$ (black line). Excess counts in the first few bins (excluded from the fit) arise from false negatives: missed clicks when the system is in $\left|\mathrm{B}\right\rangle$ produce spurious short-duration no-click intervals.
@@ C Definition of a Click
In the ideal two-level model setting of ref. [38], a click corresponds to instantaneous projection onto $\left|0\right\rangle$. In our three-level system, a click involves three sequential processes: (i) The detector drive $\Omega_{\mathrm{B}}$ induces excitations from $\left|0\right\rangle$ to $\left|\mathrm{B}\right\rangle$, (ii) the readout projects the state onto $\left|\mathrm{B}\right\rangle$, and (iii) the system relaxes back to $\left|0\right\rangle$ with relaxation time $\tau_{\mathrm{B}}$. Hence, experimentally, a “click” refers to a measurement outcome indicating detection of the system in $\left|\mathrm{B}\right\rangle$. The net effect is projection onto $\left|0\right\rangle$, as in the ideal model, but with a finite average delay $\tau_{\mathrm{B}}$. The impact of this delay on the dynamics is analysed in Section 5.
@@ D Extraction of the Measurement Rate $\alpha$
From the click record, we compute histograms of no-click durations — the time the system remains within the $\left|0\right\rangle$$\left|1\right\rangle$ manifold before a click occurs. With the system drive off ($\Omega_{\mathrm{S}}=0$) and the system initialised in $\left|0\right\rangle$, these distributions decay exponentially at rate $\alpha(\Omega_{\mathrm{B}})$, consistent with a Poisson process (Fig. S2d). While fitting this decay provides an initial estimate of $\alpha$, this approach is sensitive to misassignments. To obtain accurate values, we analyse the click record using a three-state hidden Markov model (HMM) [2] that accounts for misassignments and finite detection fidelity. This HMM is defined by a transition matrix $\mathbf{T}$ and an emission matrix $\mathbf{E}$. The transition matrix encodes the probabilities $p_{ij}$ of moving from state $\left|i\right\rangle$ to state $\left|j\right\rangle$ in one time step $dt$. In our system, transitions occur between $\left|0\right\rangle$ and $\left|\mathrm{B}\right\rangle$ and between $\left|0\right\rangle$ and $\left|1\right\rangle$, but not directly between $\left|\mathrm{B}\right\rangle$ and $\left|1\right\rangle$. The matrix therefore has the form:
@@
Figure S3: Extraction of measurement parameters via HMM analysis. a, Measurement rate $\alpha=\gamma_{\mathrm{B}\uparrow}$ versus detector drive amplitude $\Omega_{\mathrm{B}}$. While the measurement rate initially follows the expected quadratic scaling (solid black line), saturation occurs at high drive strengths ($\lambda\gtrsim 2$). Dashed grey lines indicate the observed transition values $\lambda_{1}^{\mathrm{obs}}=0.99$ and $\lambda_{2}^{\mathrm{obs}}=0.92$. Inset: expanded view of the region near the dynamical transitions. b, Decay rate $\gamma_{\mathrm{B}\downarrow}=\tau_{\mathrm{B}}^{-1}$ versus $\Omega_{\mathrm{B}}$. $\gamma_{\mathrm{B}\downarrow}$ is less affected by the drive owing to the Stark shift induced by readout photons. Error bars represent $1\sigma$ confidence intervals.
We use the Baum-Welch algorithm [2] to find the parameters that maximise the likelihood of the observed click record. Initial guesses for $\gamma_{1\uparrow}$ and $\gamma_{1\downarrow}$ come from independent $T_{1}$ and thermal population measurements, those for $\gamma_{\mathrm{B}\uparrow}$ and $\gamma_{\mathrm{B}\downarrow}$ from exponential fits to click/no-click intervals, and those for $p_{\mathrm{FP}}$ and $p_{\mathrm{FN}}$ from the Gaussian separation in the readout histogram. From the optimised transition matrix, we extract $\alpha$ and $\tau_{\mathrm{B}}$, which are used in all numerical simulations (Supplementary Information Sections 5 and 8).
@@
SI 4 Overview of the Ideal Model of Monitored Qubit Dynamics
@@ A No-Click Dynamics of a Monitored Qubit
Since the dynamics are constrained to the $XZ$-plane, the qubit state is fully characterised by the polar angle $\theta(t)$ of its Bloch vector $\left|\psi[\theta(t)]\right\rangle=\cos\!\left(\frac{\theta(t)}{2}\right)\left|1\right\rangle+\sin\!\left(\frac{\theta(t)}{2}\right)\left|0\right\rangle$. Note that $\theta=0$ corresponds to the excited state $|1\rangle$ and $\theta=\pi$ to the ground state $|0\rangle$, opposite to the convention in ref. [38]. The combined evolution leads to the equation of motion:
@@ C Angular Distributions
The second transition is characterised in ref. [38] through the steady-state angular probability density $P_{\infty}(\theta)$:
@@
Here, we instead characterise the second transition using the angular dwell-time density $\tau_{\theta}(\theta)$. Consider a system initialised at $\theta=\pi$. We define the expectation value $\mathbb{E}[T_{\Delta\theta}(\theta)]$ of the time $T_{\Delta\theta}(\theta)$ spent in the interval $\theta,\,\theta+\Delta\theta)$ before a click occurs. The dwell-time density is then
@@
When initialising the system at $\theta=0$ (as in the main text), rather than at $\theta=\pi$ (as above), there are two differences: (i) the system does not visit $\theta>0$ before a click occurs; (ii) $\tau_{\theta}(\theta)$ is multiplied by a constant factor because in the new setting $P^{(0)}(\theta=0)=1$ (as opposed to $P^{(0)}(\theta=\pi)=1$). This ultimately yields formula (4) used to fit the experimental data curves $\tau_{\theta}(\lambda,\theta)$ close to the fixed point $\theta_{+}$, and to identify the observed transition as $\lambda^{\mathrm{obs}}_{2}\equiv\lambda(\xi=0)$ (Fig. 4 and Methods).
SI 5 Realistic Model of Monitored Qubit Dynamics
In this section, we extend the ideal model of Supplementary Information Section 4 and ref. [38] to account for realistic experimental conditions:
  • Pure dephasing, modelled as Markovian fluctuations of the relative phase between $\left|0\right\rangle$ and $\left|1\right\rangle$, quantified by the pure dephasing time $T^{\phi}_{\mathrm{S}}$. This dephasing is primarily induced by the readout cavity drive.
@@
  • Detector inefficiency, manifesting as false positives (spurious clicks at rate $\kappa_{\mathrm{FP}}\equiv p_{\mathrm{FP}}/dt$, cf. Supplementary Information Section D) and false negatives (true clicks missed with probability $p_{\mathrm{FN}}$);
@@
  • Finite integration time $T_{\mathrm{int}}$, arising because the detection signal is integrated over a nonzero window, reducing temporal resolution compared to the idealised instantaneous measurements.
@@ B Detector Inefficiency
The eigenvalue structure of $\tilde{\mathcal{L}}_{P}$ determines the first two transition points, as analysed in Section 6.
@@
SI 6 The Transition Locations in the Realistic Model
We analyse how decoherence modifies the locations of the dynamical transitions, building on the realistic model of Section 5. Since the eigenvalue structure of the relevant Liouvillians fully determines the transitions, we compute the critical measurement strengths numerically by exact diagonalisation.

### v2

@@ Abstract
To probe this transition, we perform state tomography after variable evolution times without conditioning on the click record (Fig. 5A). For small $\lambda$, the decay rate grows monotonically with $\lambda$, while the oscillation frequency decreases (Fig. 5B). However, at the observed transition point $\lambda^{\mathrm{obs}}_{3}=1.09\pm 0.01$ (Methods), the dynamics become overdamped and oscillations vanish. Beyond this point, further increasing $\lambda$ slows down the exponential decay, completing the entry into the quantum Zeno regime [40]. The observed transition lies well below the ideal prediction $\lambda^{\mathrm{c}}_{3}=2$ owing to the finite waiting time $\tau_{\mathrm{B}}$ in the detector’s excited state between a click and the start of a new trajectory. This waiting time renders the dynamics partially non-Hermitian, lowering the critical value (supplementary text section D[23].
Figure 6: Summary of transition locations. Comparison of the three critical measurement strengths $\lambda^{c}_{1,2,3}$ across three levels of description: the ideal model ($\lambda^{\mathrm{c}}_{1}=1$, $\lambda^{\mathrm{c}}_{2}=2/\sqrt{3}$, $\lambda^{\mathrm{c}}_{3}=2$), the realistic model (supplementary text sections 5 and 6) incorporating decoherence ($\lambda^{\mathrm{c}}_{1}=1.18$, $\lambda^{\mathrm{c}}_{2}=1.01$, $\lambda^{\mathrm{c}}_{3}=1.25$), and experimental observations ($\lambda^{\rm obs}_{1}=0.99$, $\lambda^{\rm obs}_{2}=0.92$, $\lambda^{\rm obs}_{3}=1.09$). Notably, decoherence inverts the ordering of the first two transitions ($\lambda^{\mathrm{c}}_{2}<\lambda^{\mathrm{c}}_{1}$), a feature captured by the realistic model. Remaining discrepancies likely reflect residual systematics in calibration and data analysis (supplementary text section 7).
@@ Conclusions
The transition from drive-dominated coherent evolution to measurement-dominated behavior manifests not as a gradual crossover, but as a structured cascade of sharp dynamical transitions. By accessing the full measurement record, we resolve these transitions individually, each revealing a distinct hallmark of monitored quantum systems that ensemble averages obscure. Strikingly, decoherence does not blur these transitions but fundamentally reorganizes the dynamical phase diagram (Fig. 6), inverting the ordering of the first two critical points: dephasing shifts the state-freezing threshold below the onset of continuous quantum jumps, while the detector’s finite waiting time lowers the onset of quantum Zeno dynamics (supplementary text section 6).
The structure uncovered here points to a broader landscape. Beyond measurement strength, the temporal spacing of observations has been predicted to drive its own set of transitions—between ergodic, fractal, and localized phases [33]. Additional probes, such as click counting fields [19] or trajectory-resolved entropy production [15], may uncover further structure, including topologically distinct phases [19, 31]. More broadly, the cascaded emergence of measurement-induced dynamics should extend beyond the single-qubit setting, with implications for measurement-enhanced entanglement generation [21] and measurement-induced phase transitions in many-body systems [38, 20, 41, 11].
@@ Acknowledgments
B.G. and D.G. designed and performed the experiments. B.G., D.G., and N.N. analyzed data and performed numerical simulations. L.M.J. performed the device fabrication. U.G. and S.H. developed code for the experiments. P.K., A.R., K.S., and Y.G. developed the theory. B.G., D.G., N.N., K.S., A.R., P.K., Y.G., and S.R. wrote the manuscript, with input from all authors. Y.G., F.L., K.S., and S.R. conceived and supervised the project.
@@
Supplementary Text
SM 1 Experimental System
@@
Figure S1: Experimental implementation of a qubit with tunable monitoring strength. (A) Two-transmon chip mounted within a rectangular 3D superconducting cavity. In the expanded view, the detector (“bright”) transmon (red) is aligned with the cavity field lines (yellow arrows) to ensure strong dispersive coupling. The system (“dark”) transmon (blue) is oriented perpendicular to the field, strongly suppressing its dispersive coupling to the readout mode. (B) Energy level diagram showing the shared ground state $\left|0\right\rangle$ and two single-excitation states $\left|1\right\rangle$ and $\left|\mathrm{B}\right\rangle$ forming a V-type structure. Higher-excitation states (e.g., the doubly excited state, shown in gray) lie outside the operating manifold. A resonant drive of amplitude $\Omega_{\mathrm{B}}$ couples $\left|0\right\rangle$ to $\left|\mathrm{B}\right\rangle$, whose population is continuously monitored via the readout cavity at rate $\Gamma$. (C) This configuration effectively realizes a two-level system whose ground state is continuously monitored by a click detector with tunable measurement strength $\alpha\propto\Omega_{\mathrm{B}}^{2}/\Gamma$.
SM 2 System Parameters
@@ A Drive Configuration
To maximize detection fidelity for $\left|\mathrm{B}\right\rangle$ while minimizing decoherence in the system manifold $\{\left|0\right\rangle,\left|1\right\rangle\}$, we drive the readout at frequency $\omega_{\mathrm{R}}-\chi_{\mathrm{B}}$, which is resonant with the cavity only when the system is in $\left|\mathrm{B}\right\rangle$. The resulting monitoring rate of the $\left|\mathrm{B}\right\rangle$ population is [4]
@@ B Continuous Monitoring Protocol
Figure S2: Click record extraction. (A) Raw integrated measurement signal versus time. (B) Histogram of measurement outcomes, fitted with a sum of two Gaussians, with typical false-negative/false-positive probabilities of $p_{\mathrm{FP}}=p_{\mathrm{FN}}=4\%$. The dashed line indicates the discrimination threshold; the gray band marks the uncertainty region used for Schmitt-trigger filtering. (C) Resulting binary click record. (D) Distribution of no-click durations with $\Omega_{\mathrm{S}}=0$, showing exponential decay at rate $\alpha$ (black line). Excess counts in the first few bins (excluded from the fit) arise from false negatives: missed clicks when the system is in $\left|\mathrm{B}\right\rangle$ produce spurious short-duration no-click intervals.
@@ C Definition of a Click
In the ideal two-level model setting of ref. [40], a click corresponds to instantaneous projection onto $\left|0\right\rangle$. In our three-level system, a click involves three sequential processes: (i) The detector drive $\Omega_{\mathrm{B}}$ induces excitations from $\left|0\right\rangle$ to $\left|\mathrm{B}\right\rangle$, (ii) the readout projects the state onto $\left|\mathrm{B}\right\rangle$, and (iii) the system relaxes back to $\left|0\right\rangle$ with relaxation time $\tau_{\mathrm{B}}$. Hence, experimentally, a “click” refers to a measurement outcome indicating detection of the system in $\left|\mathrm{B}\right\rangle$. The net effect is projection onto $\left|0\right\rangle$, as in the ideal model, but with a finite average delay $\tau_{\mathrm{B}}$. The impact of this delay on the dynamics is analyzed in Section 5.
@@ D Extraction of the Measurement Rate $\alpha$
From the click record, we compute histograms of no-click durations — the time the system remains within the $\left|0\right\rangle$$\left|1\right\rangle$ manifold before a click occurs. With the system drive off ($\Omega_{\mathrm{S}}=0$) and the system initialized in $\left|0\right\rangle$, these distributions decay exponentially at rate $\alpha(\Omega_{\mathrm{B}})$, consistent with a Poisson process (Fig. S2D). While fitting this decay provides an initial estimate of $\alpha$, this approach is sensitive to misassignments. To obtain accurate values, we analyze the click record using a three-state hidden Markov model (HMM) [2] that accounts for misassignments and finite detection fidelity. This HMM is defined by a transition matrix $\mathbf{T}$ and an emission matrix $\mathbf{E}$. The transition matrix encodes the probabilities $p_{ij}$ of moving from state $\left|i\right\rangle$ to state $\left|j\right\rangle$ in one time step $dt$. In our system, transitions occur between $\left|0\right\rangle$ and $\left|\mathrm{B}\right\rangle$ and between $\left|0\right\rangle$ and $\left|1\right\rangle$, but not directly between $\left|\mathrm{B}\right\rangle$ and $\left|1\right\rangle$. The matrix therefore has the form:
@@
Figure S3: Extraction of measurement parameters via HMM analysis. (A) Measurement rate $\alpha=\gamma_{\mathrm{B}\uparrow}$ versus detector drive amplitude $\Omega_{\mathrm{B}}$. While the measurement rate initially follows the expected quadratic scaling (solid black line), saturation occurs at high drive strengths ($\lambda\gtrsim 2$). Dashed gray lines indicate the observed transition values $\lambda_{1}^{\mathrm{obs}}=0.99$ and $\lambda_{2}^{\mathrm{obs}}=0.92$. Inset: expanded view of the region near the dynamical transitions. (B) Decay rate $\gamma_{\mathrm{B}\downarrow}=\tau_{\mathrm{B}}^{-1}$ versus $\Omega_{\mathrm{B}}$. $\gamma_{\mathrm{B}\downarrow}$ is less affected by the drive owing to the Stark shift induced by readout photons. Error bars represent $1\sigma$ confidence intervals.
We use the Baum-Welch algorithm [2] to find the parameters that maximize the likelihood of the observed click record. Initial guesses for $\gamma_{1\uparrow}$ and $\gamma_{1\downarrow}$ come from independent $T_{1}$ and thermal population measurements, those for $\gamma_{\mathrm{B}\uparrow}$ and $\gamma_{\mathrm{B}\downarrow}$ from exponential fits to click/no-click intervals, and those for $p_{\mathrm{FP}}$ and $p_{\mathrm{FN}}$ from the Gaussian separation in the readout histogram. From the optimized transition matrix, we extract $\alpha$ and $\tau_{\mathrm{B}}$, which are used in all numerical simulations (supplementary text sections 5 and 8).
@@
SM 4 Overview of the Ideal Model of Monitored Qubit Dynamics
@@ A No-Click Dynamics of a Monitored Qubit
Since the dynamics are constrained to the $XZ$-plane, the qubit state is fully characterized by the polar angle $\theta(t)$ of its Bloch vector $\left|\psi[\theta(t)]\right\rangle=\cos\!\left(\frac{\theta(t)}{2}\right)\left|1\right\rangle+\sin\!\left(\frac{\theta(t)}{2}\right)\left|0\right\rangle$. Note that $\theta=0$ corresponds to the excited state $|1\rangle$ and $\theta=\pi$ to the ground state $|0\rangle$, opposite to the convention in ref. [40]. The combined evolution leads to the equation of motion:
@@ C Angular Distributions
The second transition is characterized in ref. [40] through the steady-state angular probability density $P_{\infty}(\theta)$:
@@
Here, we instead characterize the second transition using the angular dwell-time density $\tau_{\theta}(\theta)$. Consider a system initialized at $\theta=\pi$. We define the expectation value $\mathbb{E}[T_{\Delta\theta}(\theta)]$ of the time $T_{\Delta\theta}(\theta)$ spent in the interval $\theta,\,\theta+\Delta\theta)$ before a click occurs. The dwell-time density is then
@@
When initializing the system at $\theta=0$ (as in the main text), rather than at $\theta=\pi$ (as above), there are two differences: (i) the system does not visit $\theta>0$ before a click occurs; (ii) $\tau_{\theta}(\theta)$ is multiplied by a constant factor because in the new setting $P^{(0)}(\theta=0)=1$ (as opposed to $P^{(0)}(\theta=\pi)=1$). This ultimately yields formula (S2) used to fit the experimental data curves $\tau_{\theta}(\lambda,\theta)$ close to the fixed point $\theta_{+}$, and to identify the observed transition as $\lambda^{\mathrm{obs}}_{2}\equiv\lambda(\xi=0)$ (Fig. 4 and Methods).
SM 5 Realistic Model of Monitored Qubit Dynamics
In this section, we extend the ideal model of supplementary text section 4 and ref. [40] to account for realistic experimental conditions:
  • Pure dephasing, modeled as Markovian fluctuations of the relative phase between $\left|0\right\rangle$ and $\left|1\right\rangle$, quantified by the pure dephasing time $T^{\phi}_{\mathrm{S}}$. This dephasing is primarily induced by the readout cavity drive.
@@
  • Detector inefficiency, manifesting as false positives (spurious clicks at rate $\kappa_{\mathrm{FP}}\equiv p_{\mathrm{FP}}/dt$, cf. supplementary text section D) and false negatives (true clicks missed with probability $p_{\mathrm{FN}}$);
@@
  • Finite integration time $T_{\mathrm{int}}$, arising because the detection signal is integrated over a nonzero window, reducing temporal resolution compared to the idealized instantaneous measurements.
@@ B Detector Inefficiency
The eigenvalue structure of $\tilde{\mathcal{L}}_{P}$ determines the first two transition points, as analyzed in Section 6.
@@
SM 6 The Transition Locations in the Realistic Model
We analyze how decoherence modifies the locations of the dynamical transitions, building on the realistic model of Section 5. Since the eigenvalue structure of the relevant Liouvillians fully determines the transitions, we compute the critical measurement strengths numerically by exact diagonalization.